The length of a rectangle is 6 cm less than twice the width. The area of the rectangle is 108 cm2. Find the length and width of the rectangle.
Answer by unreal4evr(11)

A = L * W
108 cm2 = L * W
L = 2W - 6 cm
108 cm2 = (2W - 6 cm) * W
108 cm2 = (2W^2 - 6W)
2W^2 - 6W - 108 cm2 = 0
(W + 6)(2W - 18)
W = -6 cm or W = 9 cm (use 9)
L = 2(9 cm) - 6 cm
L = 12 cm


//Check
108 cm2 = 12 cm * 9 cm
108 cm2 = 108 cm2


L = 12 cm W = 9 cm